Grade 128 grade maths

Factorise the following expressions.
p^2 +6p+8 ??????????????????

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

p^2+6p+8
We observed that, 8 = 4×2 and 4+2 = 6
p^2+6p+8 can be written as p2+2p+4p+8
Taking Common terms, we get
p^2+6p+8
= p^2+2p+4p+8
= p(p+2)+4(p+2)
Again p+2 is common in both the terms. = (p+2)(p+4)
This implies: p^2+6p+8 = (p+2)(p+4)

Thanks